# Factuality of Large Language Models in the Year 2024

### v2

@@ Abstract
Large language models (LLMs), especially when instruction-tuned for chat, have become part of our daily lives, freeing people from the process of searching, extracting, and integrating information from multiple sources by offering a straightforward answer to a variety of questions in a single place. Unfortunately, in many cases, LLM responses are factually incorrect, which limits their applicability in real-world scenarios. As a result, research on evaluating and improving the factuality of LLMs has attracted a lot of research attention recently. In this survey, we critically analyze existing work with the aim to identify the major challenges and their associated causes, pointing out to potential solutions for improving the factuality of LLMs, and analyzing the obstacles to automated factuality evaluation for open-ended text generation. We further offer an outlook on where future research should go.
@@ 1 Introduction
To this end, researchers have collected multiple datasets, introduced a variety of measures to evaluate the factuality of LLMs, and proposed numerous strategies leveraging external knowledge through retrieval, self-reflection, and early refinement in model generation to mitigate factual errors Tonmoy et al. (2024). A number of surveys, discussing LLM hallucinations, have also been published, which we summarize in Table 1.
Why this survey?
Survey Date Pages
Eval-uation
Improve-ment
Multi-modal
Contributions and limitations
This work 01-Feb-2024 9
@@
Discusses ambiguous concepts in LLM factuality, compares and analyzes evaluation and enhancement approaches from academic and practical perspectives, outlining major challenges and promising avenues to explore.
Tonmoy et al. (2024) 08-Jan-2024 19
@@
Summarizes recent work in terms of mitigating LLM hallucinations, but lacks comparison between different approaches and discussions to identify open questions and challenges.
Gao et al. (2023b) 18-Dec-2023 26
@@
Summarizes three RAG paradigms: naïve, advanced, and modular RAG, with key elements and evaluation methods for the three major components in RAG (retriever, generator, and augmentation).
Huang et al. (2023) 09-Nov-2023 49
@@
Analyzes the reasons for hallucinations, and presents a comprehensive overview of hallucination detection methods, benchmarks, and approaches to mitigate hallucinations.
Wang et al. (2023b) 18-Oct-2023 44
@@
Detailed literature review of factuality improvement and enhancement methods covering both retrieval augmentation and non-retrieval augmentation, missing discussion of major bottleneck issues in LLM factuality and promising directions to investigate.
Rawte et al. (2023b) 18-Sept-2023 11
@@
Extensively elucidates the problem of hallucination across all major modalities of foundation models, including text (general, multilingual, domain-specific LLMs), image, video, and audio. However, inadequate coverage of approaches, in-depth categorization and comparison between methods.
Zhang et al. (2023c) 03-Sept-2023 32
@@
Organized by different training stages of LLMs, discusses potential sources of LLM hallucinations and in-depth review of recent work on addressing the problem.
Guo et al. (2022) Feb-2022 29
Focused on the automated fact-checking pipeline
@@
Table 1: Comparison of different surveys on the factuality of LLMs.
We can see in Table 1 that the majority of recent surveys on the factuality of LLMs or hallucination evaluation and improvements are rather lengthy, with most over 20 pages. While they attempt to provide a comprehensive literature review by summarizing interesting ideas and findings, organized by their intended use at different stages in the training cycle of an LLM covering pre-training, supervised fine-tuning (SFT), reinforcement learning with human or automatic feedback (RLXF), and inference, it is difficult for readers to understand the differences between some ambiguous concepts (e.g., LLM factuality vs. hallucination), fundamental challenges and promising solutions in factuality evaluation and enhancement. This survey, instead, focuses on recent novel and representative works in each category, summarizes their common bottlenecks and solutions to mitigate such issues while highlighting our personal stance based on practical experience and observations.
In addition to LLM hallucination detection and mitigation strategies, Huang et al. (2023); Zhang et al. (2023c) analyze potential reasons for LLM hallucination. Gao et al. (2023b) focus on retrieval augmentation generation (RAG) techniques whereas Guo et al. (2022) target automated fact-checking systems. The work of Tonmoy et al. (2024) compiles many recent research studies up to November 2023 albeit lacking an in-depth discussion. Rawte et al. (2023b) review numerous efforts exploring hallucinations in foundational models across various modalities, offering an overview of considerable scope and breadth, but with limited depth. Our objective is to address these challenges by concisely providing comprehensive insights into LLM factuality, prioritizing recent studies to accommodate the rapidly evolving landscape.
Scope
We outline the scope of this survey below:
  • Unlike most prior surveys, we discuss factuality across a number of modalities, including vision and speech, with an emphasis on factual errors in text.
  • We focus on factuality of open-ended text generation without references, as opposed to generation tasks with references (e.g., summarization or dialogue with source documents or user-input contexts).
  • We pay particular attention to factual errors involving world knowledge, and we leave other types (e.g., coding, reasoning, conflicting with user input or previously generated contexts) for future work.
  • In addition to improving factuality and correcting factual errors, another line of research investigates improving LLM calibration, making models aware of their limitations in its knowledge and confidence in its output, enabling it to either reject to respond to instructions or express uncertainties in its answers when unsure. Given the in-depth discussion in Geng et al. (2023), We exclude in this survey.
Structure
We first compare and contrast factuality in the context of LLMs against three closely related concepts: hallucination, relevance and trustworthiness in §2. Next, we outline benchmarks and evaluation metrics based on data format in §3. We then discuss recent works on addressing factual errors in §4 for text and other modalities in §5. Lastly, we highlight open questions and their promising solutions in §6.
@@ 2 Background
In the age of LLMs, the term hallucination has been reimagined, encompassing any deviation from factual reality or the inclusion of fabricated elements within generated texts Tonmoy et al. (2024); Rawte et al. (2023b). Zhang et al. (2023c) define hallucination as the characteristic of LLMs to generate content that diverges from the user input, contradicts previously generated context, or mis-aligns with established world knowledge. Huang et al. (2023) merge the input- and context-conflicting types of hallucinations and further take logical inconsistency into account to form faithfulness hallucination. Another category is factuality hallucination, referring to the discrepancy between generated content and verifiable real-world facts, manifesting as (1) factual inconsistency and (2) factual fabrication.
@@
Summary: Factuality is the ability of LLMs to generate content consistent with factual information and world knowledge. Although both hallucinations and factuality issues may impact the credibility of LLMs in the context of content generation, they present distinct challenges. Hallucinations occur when LLMs produce baseless or untruthful content, not grounded in the given source. In contrast, factuality errors arise when the model fails to accurately learn and utilize factual knowledge. It is possible for a model to be factually correct yet still produce hallucinations by generating content that is either off-topic or more detailed than what is requested.
Relevance vs. Factuality
FELM Chen et al. (2023) categorizes factual errors into four groups to better understand and identify LLM vulnerabilities: knowledge, irrelevance, reasoning and math, and misunderstanding falsehoods and jokes in prompts. Irrelevant error refers to that the generated content is unrelated to the prompt. For example, if the prompt is What’s a country where most people love playing rugby? a response like New Zealand is a country where rugby is considered a national passion and is deeply ingrained in the culture … would be labeled as irrelevant though it is factually true. While not factually incorrect, this response does not provide much helpful information.
@@
In the context of LLMs, factuality Wang et al. (2023b) refers to a model’s capability of generating contents of factual information, grouneded in reliable sources (e.g., dictionaries, Wikipedia or textbooks), with commonsense, world and domain-specific knowledge taken into account. In contrast, “trustworthiness” Sun et al. (2024) extends beyond mere factual accuracy and is measured on six dimensions: truthfulness, safety, fairness, robustness, privacy, and ethics.
@@ 3.1 Datasets and Metrics
Current benchmarks largely assess factuality in LLMs based on two primary capabilities: proficiency in distinguishing factual accuracy within context and ability to generate factually sound content.
@@
Perhaps the most demanding, yet inherently realistic scenario is free-form long text generation, such as biography generation. For such generations, the most commonly used and reliable methods rely on human experts following specific guidelines, and automatic fact-checkers based on retrieved information, such as FactScore, Factool and Factcheck-GPT, to facilitate efficient and consistent evaluation.
@@ 3.2 Other Metrics
In addition to evaluation methods discussed above, Lee et al. (2022) quantify hallucinations using two metrics, both requiring document-level ground-truth: (1) hallucinated named entities error measures the percentage of named entities in generations that do not appear in the ground-truth Wikipedia document; (2) entailment ratio evaluates the number of generations that can be entailed by the ground-truth reference, against the number of all generations.
Rawte et al. (2023a) define hallucination vulnerability index (HVI) to evaluate and rank LLMs based on their vulnerability to producing hallucinations, which takes a spectrum of factors into account.
Some factuality measurement tasks, such as claim extraction and evidence retrieval are non-trivial to be resolved automatically. Rawte et al. (2023a) curate publicly available LLM hallucination mitigation benchmark, where LLM generations are scored by humans when automated external knowledge retrieval fails to resolve a claim clearly. While widely used for factuality evaluation, this hybrid approach (i.e., retrieval + human) may suffer from human annotation bias.
@@ 4.2 Tuning and RLXF
Sycophancy Sharma et al. (2023), known as another source of factuality errors, often arises from misalignments during SFT and RLHFOuyang et al. (2022). This problem is partially attributed to human annotators’ tendency to award higher scores to responses they like rather than those that are factually accurate. Wei et al. (2023) explore the correlation of sycophancy with model scaling and instruction tuning. They propose a synthetic-data intervention method, using various NLP tasks to teach models that truthfulness is independent of user opinions. However, one limitation is that the generalizability of their approach remains unclear for varied prompt formats and diverse user opinions.
@@ 4.3.2 ICL and Self-reasoning
Take-away: Zheng et al. (2023) evaluate the effectiveness of knowledge editing on subject-relation-object triplets, an unrealistic setting compared to open-ended free-form text assessment. we seek answers to two research questions: (1) What types of facts and to what extent can facts be edited effectively, learned by LLMs through ICL? (2) Would SFT do a better job at learning from examples that are difficult for ICL? More broadly, what is the best way to insert new facts or edit false knowledge stored in LLMs. The community may also benefit from an in-depth comparative analysis of the effectiveness of improving factuality between SFT and ICL (perhaps also RLXF).
@@
Striking a balance between preserving creative elements and avoiding the propagation of potential errors, EVER Kang et al. (2023) and “a stitch in time saves nine” Varshney et al. (2023) actively detect and correct factual errors during generation sentence by sentence. The former leverages retrieved evidence for verification, and the latter further incorporates the probability of dominant concepts in detection. Their findings suggest that timely correcting errors during generation can prevent snowballing and further improve factuality. Nonetheless, the primary concern for this iterative process of generate-verify-correct in real-time systems is latency, making it difficult to meet the high-throughput and responsiveness demand Kang et al. (2023).
@@ 4.4 Automatic Fact Checkers
The input to a claim processor is a document to check and its expected output is a list of atomic check-worthy claims or atomic verifiable facts. There is no consensus on the granularity of “atomic claims”, making consistent decomposition difficult in a practical setting. Additionally, the concept of check-worthy and verifiable claims are subjective. Consequently, the definition of an atomic check-worthy claim remains a highly debatable topic.
This naturally leads to different “gold” human-annotated atomic claims annotated following various guidelines and distinct implementation approaches to decompose a document.
Given a document, even if assuming a ground-truth list of atomic claims, it remains an open question how to assess the quality of automatically derived decomposition results. Wang et al. (2023c) start by assessing the agreement in the number of claims between the ground truth and predictions, followed by examining the semantic similarity between two claims at the same index when the claim count aligns. Entailment ratio presented in Section 3.2 is also applicable Lee et al. (2022).
@@
The only step where we can confidently evaluate its quality is the accuracy of verification, a simple binary true/false label given a document/claim. In conclusion, perhaps the most significant hurdle limiting the development and improvement of automatic fact-checkers lies in the automated assessment and quantification of quality at intermediate stages.
@@ 5 Factuality of Multimodal LLMs
More recently, multimodal LLMs (MLLMs), capable of processing visual and audio contexts (in addition to textual input), have become a new research focus. Despite impressive results, MLLMs also hallucinate and existing research primarily focuses on addressing grounding errors.
For tuning-based approaches, Sun et al. (2023) apply RLHF to vision-language alignment tasks. They introduce factually-augmented RLHF, which augments a reward model with additional factual information, such as image captions and ground-truth options, to alleviate reward hacking. Zhao et al. (2023) introduce hallucination-aware DPO by first constructing a hallucination-aware dataset consisting of style-consistent pairs with positive and negative samples curated automatically with the help of GPT-4, and then, performing DPO training with MLLMs. While these methods enhance truthfulness, challenges remain, in the form of expensive and intricate data generation pipelines needed for creating tuning datasets.
Another line of research focuses on post-hoc approaches. Yin et al. (2023a) introduce a training-free pipeline of five steps: concept extraction, question formulation, visual validation, claim generation, and hallucination correction. They utilize textual LLMs, object detection, and visual QA (VQA) models to fix grounding errors in MLLM responses, framing detected objects and attributes as evidence. Rawte et al. (2023b) indicate that adapting textual factuality methods to MLLMs often requires conversion from various media to text (e.g. captioning). However, with increased complexity in multimodal contents, despite the application of various techniques such as VQA, OCR and ASR, fact-checking pipelines for images, audio, and video and text are largely similar. Challenges remain in accurate claim extraction and evidence retrieval across different modalities Akhtar et al. (2023).

### v3

# Factuality of Large Language Models: A Survey

@@ Abstract
Large language models (LLMs), especially when instruction-tuned for chat, have become part of our daily lives, freeing people from the process of searching, extracting, and integrating information from multiple sources by offering a straightforward answer to a variety of questions in a single place. Unfortunately, in many cases, LLM responses are factually incorrect, which limits their applicability in real-world scenarios. As a result, research on evaluating and improving the factuality of LLMs has attracted a lot of attention recently. In this survey, we critically analyze existing work with the aim to identify the major challenges and their associated causes, pointing out to potential solutions for improving the factuality of LLMs, and analyzing the obstacles to automated factuality evaluation for open-ended text generation. We further offer an outlook on where future research should go.
Factuality of Large Language Models: A Survey
Yuxia Wang1, Minghan Wang2, Muhammad Arslan Manzoor1, Fei Liu3, Georgi Georgiev4, Rocktim Jyoti Das1, Preslav Nakov1 1MBZUAI,  2Monash University,  3Google,  4Sofia University {yuxia.wang, preslav.nakov}@mbzuai.ac.ae
@@ 1 Introduction
To this end, researchers have collected multiple datasets, introduced a variety of measures to evaluate the factuality of LLMs, and proposed numerous strategies leveraging external knowledge through retrieval, self-reflection, and early refinement in model generation to mitigate factual errors Tonmoy et al. (2024). Numerous surveys (Tonmoy et al., 2024; Huang et al., 2023a; Wang et al., 2023b) have explored factuality or hallucinations in large language models across various modalities. While they either lack in-depth discussion or are too specific to grasp the fundamental challenges, promising solutions in factuality evaluation and enhancement, and some ambiguous concepts in LLM factuality. We summarized these surveys in Table 1.
Our survey aims to bridge this gap by providing an in-depth analysis of LLM factuality, with an emphasis on recent studies to reflect the rapidly evolving nature of the field. We offer a comprehensive overview of different categorizations, evaluation methods, and mitigation techniques for LLM factuality in both language and vision modalities. Additionally, we explore a novel research avenue that seeks to improve LLM calibration. This includes making models aware of their knowledge limitations and enhancing the reliability of their output confidence.
Survey Date Pages Eval Improve Multimodal Contributions and limitations
@@
Our work 15-June-2024 9 Discusses ambiguous concepts in LLM factuality, compares and analyzes evaluation and enhancement approaches from academic and practical perspectives, outlining major challenges and promising avenues to explore.
@@
Tonmoy et al. (2024) 08-Jan-2024 19 Summarizes recent work in terms of mitigating LLM hallucinations, but lacks comparison between different approaches and discussions to identify open questions and challenges.
@@
Gao et al. (2023b) 18-Dec-2023 26 Summarizes three RAG paradigms: naïve, advanced, and modular RAG, with key elements and evaluation methods for the three major components in RAG (retriever, generator, and augmentation).
@@
Huang et al. (2023b) 09-Nov-2023 49 Analyzes the reasons for hallucinations, and presents a comprehensive overview of hallucination detection methods, benchmarks, and approaches to mitigate hallucinations.
@@
Wang et al. (2023b) 18-Oct-2023 44 Detailed literature review of factuality improvement and enhancement methods covering both retrieval augmentation and non-retrieval augmentation, missing discussion of major bottleneck issues in LLM factuality and promising directions to investigate.
@@
Rawte et al. (2023b) 18-Sept-2023 11 Extensively elucidates the problem of hallucination across all major modalities of foundation models, including text (general, multilingual, domain-specific LLMs), image, video, and audio. However, inadequate coverage of approaches, in-depth categorization and comparison between methods.
@@
Zhang et al. (2023c) 03-Sept-2023 32 Organized by different training stages of LLMs, discusses potential sources of LLM hallucinations and in-depth review of recent work on addressing the problem.
Guo et al. (2022) Feb-2022 29 Focused on the automated fact-checking pipeline
@@
Table 1: Comparison of different surveys on the factuality of LLMs. Eval: Evaluation; Improve: Improvement.
@@ 2 Background
In the age of LLMs, the term hallucination has been reimagined, encompassing any deviation from factual reality or the inclusion of fabricated elements within generated texts Tonmoy et al. (2024); Rawte et al. (2023b). Zhang et al. (2023c) define hallucination as the characteristic of LLMs to generate content that diverges from the user input, contradicts previously generated context, or mis-aligns with established world knowledge. Huang et al. (2023b) merge the input- and context-conflicting types of hallucinations and further take logical inconsistency into account to form faithfulness hallucination. Another category is factuality hallucination, referring to the discrepancy between generated content and verifiable real-world facts, manifesting as (1) factual inconsistency and (2) factual fabrication.
@@
Summary: Factuality is the ability of LLMs to generate content consistent with factual information and world knowledge. Although both hallucinations and factuality may impact the credibility of LLMs in the context of content generation, they present distinct challenges. Hallucinations occur when LLMs produce baseless or untruthful content, not grounded in the given source. In contrast, factuality errors arise when the model fails to accurately learn and utilize factual knowledge. It is possible for a model to be factually correct yet still produce hallucinations by generating content that is either off-topic or more detailed than what is requested.
@@
In the context of LLMs, factuality Wang et al. (2023b) refers to a model’s capability of generating contents of factual information, grouneded in reliable sources (e.g., dictionaries, Wikipedia or textbooks), with commonsense, world and domain-specific knowledge taken into account. In contrast, “trustworthiness” Sun et al. (2024) extends beyond mere factual accuracy and is measured on eight dimensions: truthfulness, safety, fairness, robustness, privacy, ethics, transparency, and accountability.
@@ 3.1 Datasets and Metrics
Current benchmarks largely assess the factuality in LLMs based on two capabilities: proficiency in distinguishing factual accuracy in a context and ability to generate factually sound content.
@@
Perhaps the most demanding, yet inherently realistic scenario is free-form long text generation, such as biography generation. For this, the most commonly used and reliable methods rely on human experts following specific guidelines, and automatic fact-checkers based on retrieved information, such as FactScore, Factool and Factcheck-GPT, to facilitate efficient and consistent evaluation.
@@ 3.2 Other Metrics
In addition to evaluating the methods discussed above, Lee et al. (2022) quantified the hallucinations using two metrics, both requiring document-level ground-truth: (1) hallucinated named entities error measures the percentage of named entities in the generations that do not appear in the ground-truth document; (2) entailment ratio evaluates the number of generations that can be entailed by the ground-truth reference, over all generations.
Rawte et al. (2023a) defined the hallucination vulnerability index (HVI), which takes a spectrum of factors into account, to evaluate and rank LLMs.
Some factuality measurement tasks, such as claim extraction and evidence retrieval are non-trivial to automate. Rawte et al. (2023a) curated publicly available LLM hallucination mitigation benchmark, where LLM generations are scored by humans when automated external knowledge retrieval fails to resolve a claim clearly. While widely used for factuality evaluation, this hybrid approach may suffer from human annotation bias.
@@ 4.2 Tuning and RLXF
Sycophancy Sharma et al. (2023), another source of factuality errors, often arises from misalignments during SFT and RLHFOuyang et al. (2022). This is partially attributed to human annotators’ tendency to award higher scores to responses they like rather than those that are factually accurate. Wei et al. (2023) explore the correlation of sycophancy with model scaling and instruction tuning. They propose a synthetic-data intervention method, using various NLP tasks to teach models that truthfulness is independent of user opinions. However, one limitation is that the generalizability of their approach remains unclear for varied prompt formats and diverse user opinions.
@@ 4.3.2 ICL and Self-reasoning
Take-away: Zheng et al. (2023) evaluate the effectiveness of knowledge editing on subject-relation-object triplets, an unrealistic setting compared to open-ended free-form text assessment. Previous methods Mitchell et al. (2021); Meng et al. (2022) use finetuning over texts containing specific text to improve factuality. The relationship between SFT and ICL may also been an interesting avenue to explore. More specifically, we seek answers to two research questions: (1) What types of facts and to what extent can facts be edited effectively, learned by LLMs through ICL? (2) Would SFT do a better job at learning from examples that are difficult for ICL? More broadly, what is the best way to insert new facts or edit false knowledge stored in LLMs. The community may also benefit from an in-depth comparative analysis of the effectiveness of improving factuality between SFT and ICL (perhaps also RLXF).
@@
Striking a balance between preserving creative elements and avoiding error propagation, EVER Kang et al. (2023) and “a stitch in time saves nine” Varshney et al. (2023) actively detect and correct factual errors during generation sentence by sentence. The former leverages retrieved evidence for verification, and the latter incorporates the probability of dominant concepts in detection. Their findings suggest that timely correcting errors during generation can prevent snowballing and further improve factuality. Nonetheless, the primary concern for this iterative process of generate-verify-correct in real-time systems is latency, making it difficult to meet the high-throughput and responsiveness demand Kang et al. (2023).
@@ 4.4 Automatic Fact Checkers
The input to a claim processor is a document and the expected output is a list of atomic check-worthy claims or atomic verifiable facts. There is no consensus on the granularity of “atomic claims”, making consistent decomposition difficult. Additionally, the concept of check-worthy and verifiable claims are subjective. Consequently, the definition of an atomic check-worthy claim remains a highly debatable topic. This naturally leads to different “gold” human-annotated atomic claims annotated following various guidelines and distinct implementation approaches to decompose a document.
Given a document, even if assuming a ground-truth list of atomic claims, it is an open question how to assess the quality of automatically derived decomposition results. Wang et al. (2023c) assess the agreement in the number of claims between ground truth and predictions, followed by examining the semantic similarity between two claims at the same index when the claim count aligns. Entailment ratio presented in Section 3.2 is also applicable Lee et al. (2022).
@@
The only step where we can confidently evaluate its quality is the accuracy of verification, a simple binary true/false label given a document/claim. In conclusion, perhaps the most significant hurdle for the development and improvement of automatic fact-checkers lies in the automated assessment and quantification of the quality at intermediate stages.
@@ 5 Factuality of Multimodal LLMs
Factuality or hallucination in Multimodal Large Language Models refers to the phenomenon of generated responses being inconsistent with the image content. Current research on multimodal factuality can be further categorized into three types:
  1. 1.
    Existence Factuality: incorrectly claiming the existence of certain objects in the image.
  2. 2.
    Attribute Factuality: describing the attributes of certain objects in a wrong way, e.g. identifying the colour of a car incorrectly.
  3. 3.
    Relationship Factuality: false descriptions of relationships between objects, such as relative positions and interactions.
Evaluation
CHAIR (Rohrbach et al., 2018) is the first benchmark for assessing the accuracy of object existence within captions, focusing on a predefined set of objects in the COCO dataset (Lin et al., 2014). However, this approach can be misleading since the COCO dataset is frequently used in training sets, providing a limited perspective when used as the sole basis for evaluation. In contrast, POPE Li et al. (2023) evaluates object hallucination with multiple binary choice prompts, both positive and negative, querying if a specific object exists in the image. More recently, Li et al. (2023) proposed GPT4-Assisted Visual Instruction Evaluation (GAVIE) to evaluate the visual hallucination Additionally, Gunjal et al. (2023) demonstrated the use of human evaluation to avoid inaccuracies and systematic biases.
Mitigation
The methods for improving factuality in MLLMs can be broadly categorized into the categories: finetuning-based method, inference time correction and representation learning.
Fine-tuning methods such as LRV-Instruction Liu et al. (2023) and LLaVA-RLHF Sun et al. (2023) follow an intuitive and straightforward solution of collecting specialized data such as positive and negative instructions or human preference pairs. This data is used for finetuning the model, thus resulting in models with fewer hallucinated responses. Whereas inference time approaches mitigate factuality by correcting output generation. Woodpecker Yin et al. (2023a) and LURE Zhou et al. (2023) use specialized models to rectify model generation. There are other works such as HallE-Switch Zhai et al. (2023), VCD Leng et al. (2023), and HACL Jiang et al. (2023) that analyse and improve feature representation to improve factuality.